### v1

@@ 6.4 Period-luminosity relations
The radial fundamental and low-order p-modes probe the stellar mean density. For a given type of pulsator this translates into a dependency of the frequency on the luminosity. In earlier works on $\beta$ Cep stars several authors attempted to find a period-luminosity relation (e.g. Blaauw & Savedoff, 1953; Jakate, 1980; Waelkens, 1981) similar to the one in Cepheid variables (Leavitt & Pickering, 1912). However, the uncertain mode identifications led to a large scatter in these relations (see review by Lesh & Aizenman, 1978).
As we now possess complete mode identifications of tens of $\beta$ Cep stars, we can select the stars pulsating in modes that probe the mean density and fit a period luminosity relations to them. In Fig. 14, we show them in the $\log P$-$\log L/L_{\sun}$ plane colour coded by their mode degree. Using a linear regression, we arrive at a period luminosity relation for each set of pulsators:
$\log L/L_{\sun}$ $=(1.73\pm 0.41)\log P+(5.5\pm 0.30)\quad(l=0)$ (3)
$\log L/L_{\sun}$ $=(1.54\pm 0.39)\log P+(5.6\pm 0.29)\quad(l=1)$ (4)
$\log L/L_{\sun}$ $=(2.33\pm 0.75)\log P+(6.2\pm 0.59)\quad(l=2)$ (5)
with $P$ as the mode period in days.
@@
We find the three relations to be stacked by their mode degree, such that the lowest degree modes (radial) follow a period-luminosity relation with the lowest luminosity for a given period. The relation for the quadrupole modes intersects with the relation for the dipolar modes likely due to fewer stars in the sample and larger uncertainty on its parameters. Shi et al. (2024) recently provided a similar period-luminosity relation for their sample of $\beta$ Cep stars. In comparison to a linear regression of all 155 stars from Shi et al. (2024), regardless of their mode degree, our period-luminosity relations are slightly steeper. When including all stars of our sample, we find a similar slope, highlighting the importance of mode identification for period luminosity relations.
@@
Figure 14: Period luminosity relations for modes of different degrees ($l=0$ (green), $l=1$ (blue), and $l=2$ (orange)). The period luminosity relations include only the radial fundamental mode ($n=0$) for radial modes and p-modes for non-radial modes. The black line shows the relation from Shi et al. (2024) fitted to their whole sample. For comparison, we show a similar period-luminosity relation for our sample in purple (dashed). For clarity, we show the $1\sigma$-uncertainty band only for the radial pulsators. The period-luminosity relation for dipole p-mode pulsators has a similar uncertainty band while it is larger for the quadrupole pulsators. The small black data points in background shows dominant g-mode and higher-order radial pulsators excluded from the fits.
@@ 6.5 Average envelope rotation
Figure 15: Envelope rotation $\Omega_{\rm env}$ and derived parameters against observed frequencies and estimated core hydrogen fractions for 48 stars. The top panel shows the envelope rotation rate against the zonal mode frequency of the multiplet ($f_{0}$) and the second panel against the projected surface rotation rate $f_{\mathrm{rot}}\sin i$. The third and fourth panel display derived rotational properties against the evolutionary stage. In the former, we show an upper limit for the level of differential rotation ($\Omega_{\mathrm{env}}/\Omega_{\mathrm{surf}}$) and in the bottom panel the spin parameter ($s$) of the mode in the envelope. The colour-code corresponds to the mode degree: $l=1$, blue and $l=2$ orange.
@@
In Fig. 15, we show the estimated $\Omega_{\rm env}$ and derived parameters. Unlike the observed projected surface rotation and as expected, the envelope rotation deduced from the modelling does not have any correlation with the pulsation frequency of the zonal mode, as $f_{0}$ is unaffected by $\Omega_{\rm env}$ when treating the rotation perturbatively and only up to first order as we do here (top panel). The highest envelope rotation rate is not found for the fastest surface rotators but for intermediately-fast rotating stars (second panel), reflecting the effect of the unknown factor $\sin\,i$ and/or differential rotation between the envelope and the surface. This is visually represented in the third panel of Fig. 15, which shows the level of differential rotation assuming an almost equator-on view ($\sin\,i\simeq 1$). This lower limit of the true surface rotation frequency is denoted here as $\Omega_{\rm surf}$. It is used to estimate an upper limit for the differential rotation between envelope and surface, which is plotted against the evolutionary stage $X_{c}^{\prime}$. Aside from the projection effect, we find the most differential rotators to have an internal rate that is below four times the surface rotation. This is completely in line with the first such measurement ever done for a $\beta$ Cep star (Aerts et al., 2003), which relied on rotational splitting of both a dipole and quadrupole mode. Yet, the majority of stars reveal $\Omega_{\rm env}/\Omega_{\rm surf}\lesssim 2$. Hence, despite the few cases of strong radial differential rotation as summarised by Burssens et al. (2023), the majority of $\beta$ Cep stars have a level of envelope-to-surface rotation between 0.2 and 2.
@@
We also present the spin-parameter, $s$, of the dominant mode as a function of $X_{c}^{\prime}$. It is defined here as $s\equiv 2\Omega_{\mathrm{env}}/f_{0}$ and characterises the regime in which the pulsation mode occurs (Aerts & Tkachenko, 2024). All the Gaia $\beta$ Cep stars have their dominant mode in the super-inertial regime with $s<1$. Noticably, the star with the largest spin parameter (TIC 13765605) also has the highest level of differential rotation and is about halfway its main-sequence phase.
@@
Figure 16: Histograms of $\epsilon^{2}$ (top) giving the influence of the centrifugal force with respect to gravity and the inverse of the $K$-value (bottom) representing the importance of the 2nd-order Coriolis force versus the centrifugal force in the treatment of the pulsations.
@@
Finally, we show in Fig. 16 histograms of the dimensionless quantities $\epsilon^{2}\equiv(2\pi\Omega_{\rm env})^{2}R^{3}/GM$ and $K^{-1}=(2\pi f_{0})^{2}R^{3}/GM$, following Goupil et al. (2000). The latter quantity is the inverse of $K$ already introduced above. It represents an assessment of the importance of the centrifugal force with respect to the 2nd-order Coriolis force in the treatment of the pulsations, while the former evaluates the effect of the centrifugal force with respect to gravity. Goupil et al. (2000) reported the centrifugal force (and hence the stellar oblateness) to be more important for the computation of the p-modes of rapidly rotating $\delta\,$Sct stars compared to the 2nd-order rotational effect caused by the Coriolis force. Here, we provide the distributions of both $\epsilon^{2}$ and $K^{-1}$ for our population of Gaia $\beta$ Cep stars. While their mode frequencies are in general lower than those of $\delta\,$Sct stars, we still find the second-order Coriolis force to be less important than the centrifugal force for the computation of the dominant modes of the Gaia $\beta$ Cep stars, but the difference in importance between the two forces is smaller than for $\delta\,$Sct stars. We observe a bimodal distribution for the modes in the $\beta$ Cep population, showing that the $l=2$ modes have higher dependence on the centrifugal deformation than the $l=1$ modes.
@@ 7 Conclusions
We analysed the TESS light curves of 216 Gaia discovered $\beta$ Cep stars (Gaia Collaboration et al., 2023), including 145 new detections. By placing these stars in a Gaia HRD, we confirmed that all of the stars are late O or early B stars near the $\beta$ Cep instability region. Yet, a detailed light curve analysis shows that at least three stars are short period eclipsing binaries. By categorising the TESS light curves, we find the majority of $\beta$ Cep stars in our sample to be multi-periodic.
@@
The combination of the Gaia and TESS time series photometry, allowed us to determine the mode degrees for 176 stars in our sample from their amplitude ratios. Most observed $\beta$ Cep stars pulsate in non-radial modes with $l=1$. Our identified modes constitute the largest set of identified modes of $\beta$ Cep stars available to date.
@@
From the TESS light curves, we found 60 rotationally split multiplets accompanying the main pulsation mode. The majority of these multiplets have positive asymmetries in their splittings. Yet, several stars show negative asymmetries and are prime targets for detailed follow-up analyses. We provide the first distributions of the envelope rotation rates and of upper limits for the level of differential rotation for a meaningful sample of $\beta$ Cep stars. We find that most stars have envelope-to-surface ratios less than two and many of them rotate almost rigidly. We do point out that our sample only consists of $\beta$ Cep stars with a dominant mode in the super-inertial regime and thus lacks representation of the fastest rotating class members due to our sample selection from Gaia.
@@
Further, we model 164 stars based on the grid calculated by Burssens et al. (2023). We provide distributions in age, convective core mass, and mass in order to inform population studies of $\beta$ Cep stars. We find the $l=2$ mode pulsators to be distributed bimodal with respect to the core hydrogen fraction and the radial mode pulsators to occur mostly in the intermediate main-sequence stage. The period-luminosity relations for the different mode degrees are steeper than one obtained for the whole ensemble.
@@
With the upcoming Gaia data releases and the on-going TESS mission our approach will be applicable to the majority of known and still to be discovered $\beta$ Cep stars. With a longer time base and improved instrumental corrections, Gaia will deliver many more new $\beta$ Cep stars with future data releases. But even with the current data, expanding our work to all known $\beta$ Cep stars by searching for lower amplitude pulsations in the Gaia data is possible. Identifying, the majority of main modes in $\beta$ Cep stars will further improve our understanding of $\beta$ Cep stars from a population point-of-view while delivering the highly needed input to ensemble forward modelling. Complementary to our photometric analysis of fainter $\beta$ Cep stars, the upcoming CubeSpec mission (Raskin et al., 2018; Bowman et al., 2022) will target bright $\beta$ Cep stars with high-resolution time series spectroscopy from space to determine their pulsation modes to revive studies of massive pulsators and unlock their full asteroseismic potential.
@@ Acknowledgements.
We thank Tim Bedding, Dominic Bowman, Siemen Burssens, Zhao Guo, Luc IJspeert, Kalarickal Sreenivas, and Vincent Vanlaer for useful discussions. The research leading to these results has received funding from the KU Leuven Research Council (grant C16/18/005: PARADISE). CA and MV acknowledge financial support from the European Research Council (ERC) under the Horizon Europe programme (Synergy Grant agreement N101071505: 4D-STAR). While partially funded by the European Union, views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. JDR and CA acknowledge the Belgian Federal Science Policy Office (BELSPO) for the provision of financial support in the framework of the PRODEX Programme of the European Space Agency (ESA) This research has made use of NASA’s Astrophysics Data System Bibliographic Services and of the SIMBAD database and the VizieR catalogue access tool, operated at CDS, Strasbourg, France. This paper includes data collected by the TESS mission. Funding for the TESS mission is provided by the NASA’s Science Mission Directorate. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
@@
Appendix A Additional figures
Here, we present some additional figures that are informative to the reader. In Fig. 17, we show an example of a light curve that is affected by instrumental trends and the final detrended light curve used in our analysis. Figure 18 shows an HRD colour-coded by the light curve classification, Fig. 19 is a repetition of Fig. 8 but this time including the mode identifications based on multi-colour photometry. Finally, we colour-code the stars in the HRD in Fig. 20 by their mode identification.
@@ Appendix A Additional figures
Figure 17: Example light curve before (top) and after (bottom) the final detrending. Sections highlighted in red in the top panel have been removed because more than ten stars in our sample display similar instrumental effects at the selected times. We note that in the top panel, a downward slope is present in the first half of the sector which has been successfully removed by our detrending.

### v2

@@ 6.4 Period-luminosity relations
Radial and low-order non-radial p-modes probe the internal density of a star. For a given type of pulsator this translates into a connection between a mode’s frequency and the luminosity as established for Cepheids by Leavitt & Pickering (1912). For the $\delta\,$Sct stars, which are situated in the low-luminosity part of the classical instability strip, this also gives rise to a period-luminosity relation (Gaia Collaboration et al., 2023) but it is not so clean as the one found for Cepheids because a fraction of these p-mode pulsators have a dominant non-radial mode (Ziaali et al., 2019).
While the $\delta\,$Sct stars occupy a small region in mass and radius, this is not the case for the $\beta\,$Cep stars. Given their large spread in mass and radius along the main sequence, one does not expect a clear period-luminosity relation for this class as a whole. Nevertheless, several authors attempted to find such a relation in the past (e.g. Blaauw & Savedoff, 1953; Jakate, 1980; Waelkens, 1981). However, the uncertain mode identifications led to a large scatter in these relations (see review by Lesh & Aizenman, 1978).
Here we show the relations and their scatter occurring for our sample of $\beta\,$Cep star when using only those with identified pulsation modes. Figure 16 shows the period-luminosity relations in the $\log P$-$\log L/L_{\sun}$ plane, colour-coded by the mode degree. Using a linear regression, we arrived at a period luminosity relation for each sub-set of pulsators according to their dominant mode:
$\log L/L_{\sun}$ $=(2.21\pm 0.52)\log P+(5.8\pm 0.36)\quad(l=0)$ (5)
$\log L/L_{\sun}$ $=(1.81\pm 0.36)\log P+(5.8\pm 0.27)\quad(l=1)$ (6)
$\log L/L_{\sun}$ $=(0.94\pm 1.34)\log P+(5.1\pm 1.02)\quad(l=2)$ (7)
with $P$ the mode period in days. Figure 16 shows only the $1\sigma$-uncertainty band for the stars with a dominant fundamental radial mode. Those with a dominant dipole p-mode have a similar uncertainty band while it is larger for the quadrupole mode pulsators.
@@
We find the three relations to be stacked by their mode degree, such that the lowest degree modes (radial) follow a period-luminosity relation with the lowest luminosity for a given period. The relation for the quadrupole modes intersects with the relation for the dipolar modes due to fewer stars in the sample and larger spread in the radial order. Shi et al. (2024) recently provided a similar period-luminosity relation for their sample of $\beta$ Cep stars. In comparison to a linear regression of all 155 stars from Shi et al. (2024), regardless of their mode degree, our period-luminosity relations are steeper. When including all stars of our sample, we find a similar slope to theirs, highlighting the importance of mode identification.
@@
Figure 16: Period-luminosity relations for stars with modes of different degrees ($l=0$ (green), $l=1$ (blue), and $l=2$ (orange)). The period-luminosity relations include only the radial fundamental mode ($n=0$) for radial pulsators and p-modes for the non-radial pulsators. The black line shows the relation of the Shi et al. (2024) sample. For comparison, a similar period-luminosity relation for our sample is plotted in purple (dashed). The green shaded area shows the uncertainty band of the fundamental radial-mode pulsators. The small black data points in the background show dominant g-mode and higher-order radial pulsators excluded from the fits.
@@ 6.5 Average envelope rotation
Figure 17: Envelope rotation $\Omega_{\rm env}$ and derived parameters against observed frequencies and estimated core hydrogen fractions for 25 stars. The top panel shows the envelope rotation rate against the zonal mode frequency of the multiplet ($f_{0}$) and the second panel against the projected surface rotation rate $f_{\mathrm{rot}}\sin i$ for the stars having this quantity available. The third and fourth panels display derived rotational properties against the evolutionary stage. In the former, we show an upper limit for the level of differential rotation ($\Omega_{\mathrm{env}}/\Omega_{\mathrm{surf}}$) and in the bottom panel the spin parameter ($s$) of the mode in the envelope. The colour-code corresponds to the mode degree: $l=1$ in blue and $l=2$ in orange.
@@
In Fig. 17, we show the estimated $\Omega_{\rm env}$ and derived parameters. Unlike the observed projected surface rotation and as expected, the envelope rotation deduced from the modelling does not have any correlation with the pulsation frequency of the zonal mode, as $f_{0}$ is unaffected by $\Omega_{\rm env}$ when treating the rotation perturbatively and only up to first order as we do here (top panel). The highest envelope rotation rate is not found for the fastest surface rotators but for intermediately-fast rotating stars (second panel), reflecting the effect of the unknown factor $\sin\,i$ and/or differential rotation between the envelope and the surface. This is visually represented in the third panel of Fig. 17, which shows the level of differential rotation assuming an almost equator-on view ($\sin\,i\simeq 1$). This lower limit of the true surface rotation frequency is denoted here as $\Omega_{\rm surf}$. It is used to estimate an upper limit for the differential rotation between envelope and surface, which is plotted against the evolutionary stage $X_{c}^{\prime}$. Aside from the projection effect, we find the most differential rotators to have an internal rate that is below four times the surface rotation. This is completely in line with the first such measurement ever done for a $\beta$ Cep star (Aerts et al., 2003), which relied on rotational splitting of both a dipole and quadrupole mode. Yet, the majority of stars reveal $\Omega_{\rm env}/\Omega_{\rm surf}\lesssim 2$. Hence, despite the few cases of strong radial differential rotation as summarised by Burssens et al. (2023), the majority of $\beta$ Cep stars have a level of envelope-to-surface rotation between 0.3 and 2.
@@
We also present the spin-parameter, $s$, of the dominant mode as a function of $X_{c}^{\prime}$. It is defined here as $s\equiv 2\Omega_{\mathrm{env}}/f_{0}$ and characterises the regime in which the pulsation mode occurs (Aerts & Tkachenko, 2024). All the Gaia $\beta$ Cep stars have their dominant mode in the super-inertial regime with $s<1$. Noticeably, the star with the largest spin parameter (TIC 282881728) also has the highest level of differential rotation and is in the later half of its main-sequence phase.
@@
Figure 18: Histograms of $\epsilon^{2}$ (top) giving the influence of the centrifugal force with respect to gravity and the inverse of the $K$-value (bottom).
@@
Finally, we show in Fig. 18 histograms of the dimensionless quantities $\epsilon^{2}\equiv(2\pi\Omega_{\rm env})^{2}R^{3}/GM$ and $K^{-1}=(2\pi f_{0})^{2}R^{3}/GM$, following Goupil et al. (2000). The latter quantity is the inverse of $K$ already introduced above. It represents an assessment of the importance of the centrifugal force with respect to the 2nd-order Coriolis force in the treatment of the pulsations, while the former evaluates the effect of the centrifugal force with respect to gravity. Goupil et al. (2000) reported the centrifugal force (and hence the stellar oblateness) to be more important for the computation of the p-modes of rapidly rotating $\delta\,$Sct stars compared to the 2nd-order rotational effect caused by the Coriolis force. While the mode frequencies of $\beta$ Cep stars are in general lower than those of $\delta\,$Sct stars, we still find the second-order Coriolis force to be less important than the centrifugal force. However, the difference in importance between the two forces is smaller than for $\delta\,$Sct stars.
@@ 7 Conclusions
We analysed the TESS light curves of 216 Gaia-detected $\beta$ Cep candidates (Gaia Collaboration et al., 2023), including 145 new detections. By placing these stars in a Gaia HRD, we confirmed that all of the stars are late O or early B stars near the $\beta$ Cep instability region. Yet, a detailed light curve analysis shows that at least three stars are short period eclipsing binaries. By categorising the TESS light curves, we find the majority of $\beta$ Cep stars in our sample to be multi-periodic.
@@
The combination of the Gaia and TESS time series photometry, allowed us to determine the mode degrees for 148 stars in our sample from their amplitude ratios. Most observed $\beta$ Cep stars pulsate in non-radial modes with $l=1$. Our identified modes constitute the largest set of identified modes of $\beta$ Cep stars available to date.
@@
From the TESS light curves, we found 34 rotationally split multiplets accompanying the main pulsation mode. The majority of these multiplets have positive asymmetries in their splittings. Yet, several stars show negative asymmetries and are prime targets for detailed follow-up analyses. We provide the first distributions of the envelope rotation rates and of upper limits for the level of differential rotation for a meaningful sample of $\beta$ Cep stars. We find that most stars have envelope-to-surface ratios less than two and many of them rotate almost rigidly. We do point out that our sample only consists of $\beta$ Cep stars with a dominant mode in the super-inertial regime and thus lacks representation of the fastest rotating class members due to our sample selection from Gaia.
@@
Further, we performed the first ensemble modelling of 119 $\beta$ Cep stars based on the grid calculated by Burssens et al. (2023). We provide distributions in age, convective core mass, and mass in order to inform population studies of $\beta$ Cep stars. With the large set of identified modes, we are in a position to provide these distributions for each mode degree. As expected from instability calculations, we find the stars to be at least one third through their main sequence evolution. The radial mode pulsators in particular occur mostly in the intermediate main-sequence stage. The period-luminosity relations for the different mode degrees are steeper than one obtained for the whole ensemble.
@@
With the upcoming Gaia data releases and the on-going TESS mission our approach will be applicable to the majority of known and still to be discovered $\beta$ Cep stars. With a longer time base and improved instrumental corrections, Gaia will deliver many more new $\beta$ Cep stars with future data releases. But even with the current data, expanding our work to all known $\beta$ Cep stars by searching for lower amplitude pulsations in the Gaia data is possible. Identifying the majority of main modes in $\beta$ Cep stars will further improve our understanding of $\beta$ Cep stars from a population point-of-view while delivering the highly needed input to ensemble forward modelling. Complementary to our photometric analysis of fainter $\beta$ Cep stars, the upcoming CubeSpec mission (Raskin et al., 2018; Bowman et al., 2022) will target bright $\beta$ Cep stars with high-resolution time series spectroscopy from space to determine their pulsation modes, while the Global Asteroseismology Project is targetting mode identification from long-term multi-colour continuous ground-based observations (Shitrit & Arcavi, 2024). These ongoing and future projects will revive studies of massive pulsators and unlock their full asteroseismic potential.
@@ Acknowledgements.
We thank the referee for the detailed and critical review that helped to improve the paper. We thank Tim Bedding, Dominic Bowman, Siemen Burssens, Luc IJspeert, Kalarickal Sreenivas, and Vincent Vanlaer for useful discussions. The research leading to these results has received funding from the KU Leuven Research Council (grant C16/18/005: PARADISE). CA and MV acknowledge financial support from the European Research Council (ERC) under the Horizon Europe programme (Synergy Grant agreement N101071505: 4D-STAR). While partially funded by the European Union, views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. JDR and CA acknowledge the Belgian Federal Science Policy Office (BELSPO) for the provision of financial support in the framework of the PRODEX Programme of the European Space Agency (ESA) This research has made use of NASA’s Astrophysics Data System Bibliographic Services and of the SIMBAD database and the VizieR catalogue access tool, operated at CDS, Strasbourg, France. This paper includes data collected by the TESS mission. Funding for the TESS mission is provided by the NASA’s Science Mission Directorate. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
@@
Appendix A Detrending and special photometry
In order to use the pulsation modes of the $\beta\,$Cep stars in our sample we ensured 1) minimal contribution by contaminating stars (Sect. 2.1) and 2) a clean light curve containing as few instrumental signals as possible. We removed instrumental and long-term trends after excluding intervals which proved problematic in multiple stars. We preferred excluding these intervals over complicated detrending steps to avoid overcorrection. Further, some weak long-term trends sometimes remained and the rising or falling curves at the sector edges were not always easily corrected. Hence, we detrended blocks of data separated by a gap of at least three hours with second-order polynomials. In the example shown in Fig. 19, more than ten stars in our sample display similar instrumental effects at the selected times. We note that in the top panel, a downward slope is present in the first half of the sector which has been successfully removed by our detrending. Afterwards, we $\sigma$-clipped the light curves with a threshold of 5 $\sigma$. Finally, in 18 light curves, hand-picked time intervals or even entire sectors were removed if they displayed unusual behaviour likely caused by instrumental effects or contamination and $\sigma$-clipped them again.
For 15 stars, we used single-pixel photometry, in order to avoid contamination from nearby sources. One example is given in Fig. 20. Although the main $\beta$ Cep pulsation is clearly visible in the periodograms of nearby pixels, several additional peaks emerge for these stars too, indicating a periodic contamination from other sources. The particular case of TIC 264026227 is a member of the open cluster NGC 5606, which host other massive stars in close proximity of TIC 264026227. We note that the Gaia detected frequency does correspond to the main TESS frequency, although the light curves in nearby pixels have other periodic components that are more prominent. Hence, using a single pixel for the photometry of this star ensures the cleanest possible extraction.
@@ Appendix A Detrending and special photometry
Figure 19: Example light curve before (top) and after (bottom) the final detrending. Sections highlighted in red in the top panel have been removed.
Figure 20: Example of a single pixel TESS light curve. The target TIC 264026227, member of the open cluster NGC 5606, is contaminated by several nearby stars. Top left: The TESS FFI cut-out of the field around TIC 264026227, with the pixel used for or final light curve marked in red. The colour scale shows the flux intensity with yellow being brighter pixels. Over-plotted in each pixel is its light curve. Top right: Same as left but with the periodogram over-plotted. Bottom: Final light curve shows clear $\beta$ Cep variability after additional cleaning and detrending beyond what is shown in the top left panel.
Appendix B Additional Hertzsprung-Russell diagrams and mode identification
For convenience of the interested reader, we provide some extra HRDs with colour-coding by the light curve classification (Fig. 21, Sect. 4.1) and mode degree as identified from multi-colour photometry (Fig. 22, Sect. 5.3). Both diagrams show no obvious structure in the light of the current study.
Figure 23 shows the distribution of amplitude ratios for the set of mode degrees as determined by our probabilistic approach. Qualitatively, the diagram resembles Fig. 9. While the Gaia amplitude ratios of the non-radial pulsators overlap significantly, the TESS amplitudes for the non-radial modes are well separated. As expected, the amplitude ratios of the radial modes drop stronger over the whole wavelength range than the amplitude ratios of the non-radial modes.